# Human Action CLIPS: Detecting AI-generated Human Motion

### v1

@@
Abstract.
Full-blown AI-generated video generation continues its journey through the uncanny valley to produce content that is perceptually indistinguishable from reality. Intermixed with many exciting and creative applications are malicious applications that harm individuals, organizations, and democracies. We describe an effective and robust technique for distinguishing real from AI-generated human motion. This technique leverages a multi-modal semantic embedding, making it robust to the types of laundering that typically confound more low- to mid-level approaches. This method is evaluated against a custom-built dataset of video clips with human actions generated by seven text-to-video AI models and matching real footage.
1. Introduction
@@
Generative AI, unlike earlier model-based animation approaches, has emerged as an intriguing new genre for computer animation. While earlier versions of AI-generated video were things nightmares are made of (see, for example, “Will Smith eating spaghetti”111https://www.youtube.com/watch?v=XQr4Xklqzw8), recent advances have shown significant improvements in photo-realism and temporal consistency.
@@
We have already started to see AI-generated images weaponized in the form of child sexual abuse material, non-consensual sexual imagery, fraud, and as an accelerant to disinformation campaigns (Farid, 2022). There is no reason, therefore, to believe that AI-generated video will not follow suit.
@@
We describe a technique to detect AI-generated videos containing human motion. Our initial focus is on humans because these videos are of most concern when it comes to the harms enumerated above. Despite this focus, we will discuss why our approach is likely to generalize to other types of video content. We evaluate the efficacy of our approach on a diverse dataset of our creation consisting of real and matching (in terms of the human actions depicted) AI-generated videos from seven different generative-AI models. We demonstrate the robustness of our detection in the face of standard laundering attacks like resizing and transcoding, and evaluate its generalizability to previously unseen models.
1.1. Related Work
Reactive techniques – operating in the absence of credentials – fall into two basic approaches: (2a) learning-based, in which features that distinguish real from fake content are learned by a range of different machine-learning techniques, and (2b) artifact-based, in which a range of low-level (pixel-based) to high-level (semantic-based) features are explicitly extracted to distinguish between real and fake content (Farid, 2022).
Another recent example (Jia et al., 2024) explores the potential of multi-modal large language models (LLMs) for detecting AI-generated faces. In this approach, the authors prompt ChatGPT with an image and prompt like “Tell me if this is an AI-generated image.” This approach achieves an average accuracy of (as measured by area under the curve, AUC). Although not particularly accurate, what is intriguing about this approach is that it points to a potentially semantic-level reasoning.
Recent studies took a more direct semantic approach by leveraging a contrastive language-image pretraining (CLIP) representation (Radford et al., 2021). In (Cozzolino et al., 2024), the authors extract a CLIP embedding from real and AI-generated images and with only a linear SVM achieve detection accuracy ranging from to depending on the amount of image post-processing (cropping, resize, recompression).
@@ 1.1. Related Work
[…]
Sample frames corresponding to the prompt “a person dancing to music” from one real and seven AI-generated videos.
1.2. Contribution
Motivated by, and building upon, earlier work, we describe new techniques tailored to detect AI-generated video containing human motion. Our work contributes to this nascent literature with the:
@@ 1.2. Contribution
construction and release of a new dataset of real and AI-generated human motion.
2. Methods
We begin by describing the collection and creation of real and AI-generated video clips consisting of varying types of human motion, followed by a description of four CLIP-like, multi-modal embeddings and the underlying classification schemes used to distinguish the real from the fake.
2.1. Dataset
We generated video clips from seven text-to-video AI models: BD AnimateDiff (Lin and Yang, 2024), CogVideoX-5B (Yang et al., 2024), Lumiere222An earlier version of the Lumiere model was used which was not specifically trained or fine-tuned for generating videos of human motion. (Bar-Tal et al., 2024), RunwayML Gen3 (Gen3, [n. d.]), Stable Diffusion Txt2Img+Img2Vid (Blattmann et al., 2023), Veo333A pre-release version of the Veo model was used. (Veo, [n. d.]), and VideoPoet (Kondratyuk et al., 2023). The default generation parameters were used for each model. These AI-generated videos are minutes in length constituting video frames. See Figure 2 for representative examples.
@@ 2.1. Dataset
We also curated a set of real videos from Pexels (Pexels, [n. d.]), an open-source stock video database, matched to the human-action prompts described above. These real videos are minutes in length constituting individual video frames. These videos were matched on general content/action so that the embedding representations (described next) between real and fake would not be based on semantic differences.
The resulting dataset, which we call DeepAction, comprises eight video categories: seven AI-generated and one real. The dataset is available at https://huggingface.co/datasets/faridlab/deepaction_v1.
2.2. Embeddings
@@
The CLIP embedding model was trained on LAION-2B (Schuhmann et al., 2022). We used the smallest -D baseline model ViT-B/32444https://huggingface.co/openai/clip-vit-base-patch32.
The SigLIP embedding model was trained on WebLI (Chen et al., 2022), which is over larger than LAION-2B. Unlike CLIP’s softmax-based contrastive loss, SigLIP uses a sigmoid loss. We used the -D patch16-224 model variant555https://huggingface.co/google/siglip-base-patch16-224.
The JinaCLIP embedding model was trained on LAION-400M (Schuhmann et al., 2021) along with an additional set of text-pair datasets. While CLIP’s training optimized for text-image representation alignment, JinaCLIP was trained to jointly optimize text-image and text-text representation alignment. We used the base -D v1 model666https://huggingface.co/jinaai/jina-clip-v1.
@@
2.3. Classification
@@
A two-class SVM (Cortes and Vapnik, 1995), implemented in scikit-learn777https://scikit-learn.org/stable/modules/generated/sklearn.svm.SVC.html, is used to classify video frames as real or fake, in which all seven text-to-video models are bundled into a single ’fake’ class.
@@ 3. Results
For the CLIP, SigLIP, and JinaCLIP embeddings, our dataset is randomly split into an train/test partition. For the fine-tuned CLIP, the dataset is randomly split into a partition, where the first is used for CLIP fine-tuning, the next is used for model training, and the remaining is used for testing.
In each case, we report the mean frame- and video-level accuracy on the test set, averaged over five random train/test repetitions. Because our dataset is imbalanced, with significantly more fake than real videos, we under-sample the fake videos to create a balanced dataset. Throughout, we report accuracy as a macro-average by evenly weighting the class accuracies.
@@
A 2D PCA visualization of the real and fake finetuned-CLIP embeddings per video frame.
@@
A 2D PCA visualization of VideoPoet and each of six other text-to-video fine-tuned CLIP embeddings per video frame, showing a distinct difference across AI models.
@@ 3.1. Robustness
Shown in the left portion of Table 3 are the frame- and video-level accuracies for the CLIP embedding and a two-class linear SVM for videos with progressively lower resolution (we quantify the change in resolution as a percentage because the original videos were of varying resolution, Section 2.1). At the video level, accuracy remains relatively high for videos at or higher of their original resolution. This corresponds to an average resolution of .
@@ 3.2. Generalizability
3.3. Non-Human Motion
Having been trained on only videos containing human motion, we wondered if our two-class SVM would generalize to arbitrary text-to-video models. We evaluated our model on videos generated by Sora (Sora, [n. d.]) and RunwayML (Gen3, [n. d.]) that did not contain any humans or human-motion. With the CLIP embeddings, the model correctly classified of these videos as AI generated showing (a somewhat surprising) generalization to non-human motion and to one text-to-video model not seen in training (Sora).
@@ 3.5. CGI
Performance for SigLIP across all prompts was perfect at . That is, the SigLIP embedding for these video frames is more similar to the “AI-generated”, “unrealistic”, or “manipulated” prompts than the “authentic” or “realistic” prompts. We don’t fully understand why there is such a large discrepancy here across embedding, but it may be possible that SigLIP – trained on WebLI – was in fact exposed to CGI content.
3.6. Comparison to Related Work
Due to the lack of standardized benchmarks, comparing methods remains imperfect. Nevertheless, we provide a comparison to related work by comparing baseline accuracy as well as generalizability to unseen generative models. The latter is particularly important because true efficacy in the wild will be limited by the ability of detection models to generalize.
@@ 3.6. Comparison to Related Work
4. Discussion
When we first started to think about the problem of distinguishing real from AI-generated human motion, we thought to take a physics-based approach leveraging recent advances in 3D human-body modeling (Loper et al., 2023). We hypothesized that by extracting 3D models of the human body, we could expose implausible dynamics and kinematics in AI-generated motion. We, however, quickly ran into obstacles and found it difficult to consistently and reliably extract 3D models in a wide range of human poses and levels of occlusion.
@@ 4. Discussion
Acknowledgements.
@@
This work was done during the first author’s (MB) internship at Google. The first author thanks his manager, Tony Chen, and co-host, Catherine Zhao, for their exceptional support, mentoring, and discussions related to this work; thanks to Prakash Hariramani, Sunita Verma, and Joshy Joseph for their support within leadership; and thanks for helpful discussions to: Amanda Walker, Athula Balachandran, Dara Bahri, David Hendon, Jess Gallegos, John Choi, Karen Guo, Nicholas Carlini, Nick Pezzoti, Stanley Wang, Sven Gowal, and Tom Hume. This work was supported by a gift from YouTube and funding from the University of California Noyce Initiative.References
- (1)
- Agarwal et al. (2020) Shruti Agarwal, Hany Farid, Tarek El-Gaaly, and Ser-Nam Lim. 2020. Detecting deep-fake videos from appearance and behavior. In IEEE International Workshop on Information Forensics and Security. IEEE, 1–6.
- Bao et al. (2024) Fan Bao, Chendong Xiang, Gang Yue, Guande He, Hongzhou Zhu, Kaiwen Zheng, Min Zhao, Shilong Liu, Yaole Wang, and Jun Zhu. 2024. Vidu: A highly consistent, dynamic and skilled text-to-video generator with diffusion models. arXiv:2405.04233.
- Bar-Tal et al. (2024) Omer Bar-Tal, Hila Chefer, Omer Tov, Charles Herrmann, Roni Paiss, Shiran Zada, Ariel Ephrat, Junhwa Hur, Yuanzhen Li, Tomer Michaeli, et al. 2024. Lumiere: A space-time diffusion model for video generation. arXiv:2401.12945.
- Barrington et al. (2023) Sarah Barrington, Romit Barua, Gautham Koorma, and Hany Farid. 2023. Single and multi-speaker cloned voice detection: From perceptual to learned features. In IEEE International Workshop on Information Forensics and Security. IEEE, 1–6.
- Barrington et al. (2024) Sarah Barrington, Matyas Bohacek, and Hany Farid. 2024. DeepSpeak Dataset v1. 0. arXiv:2408.05366.
- Barrington and Farid (2024) Sarah Barrington and Hany Farid. 2024. People are poorly equipped to detect AI-powered voice clones. arXiv:2410.03791.
- Blattmann et al. (2023) Andreas Blattmann, Tim Dockhorn, Sumith Kulal, Daniel Mendelevitch, Maciej Kilian, Dominik Lorenz, Yam Levi, Zion English, Vikram Voleti, Adam Letts, et al. 2023. Stable video diffusion: Scaling latent video diffusion models to large datasets. arXiv:2311.15127.
- Blue et al. (2022) Logan Blue, Kevin Warren, Hadi Abdullah, Cassidy Gibson, Luis Vargas, Jessica O’Dell, Kevin Butler, and Patrick Traynor. 2022. Who are you (I really wanna know)? Detecting audio DeepFakes through vocal tract reconstruction. In 31st USENIX Security Symposium. 2691–2708.
- Boháček and Farid (2022) Matyáš Boháček and Hany Farid. 2022. Protecting world leaders against deep fakes using facial, gestural, and vocal mannerisms. Proceedings of the National Academy of Sciences 119, 48 (2022), e2216035119.
- Bohacek and Farid (2024) Matyas Bohacek and Hany Farid. 2024. Lost in Translation: Lip-Sync Deepfake Detection from Audio-Video Mismatch. In IEEE/CVF Conference on Computer Vision and Pattern Recognition, Workshop on Media Forensics. 4315–4323.
- Cai et al. (2021) Zhongang Cai, Mingyuan Zhang, Jiawei Ren, Chen Wei, Daxuan Ren, Zhengyu Lin, Haiyu Zhao, Lei Yang, Chen Change Loy, and Ziwei Liu. 2021. Playing for 3D human recovery. arXiv:2110.07588.
- Carlini and Farid (2020) Nicholas Carlini and Hany Farid. 2020. Evading deepfake-image detectors with white-and black-box attacks. In International Conference on Computer Vision and Pattern Recognition Workshop. 658–659.
- Chen et al. (2022) Xi Chen, Xiao Wang, Soravit Changpinyo, AJ Piergiovanni, Piotr Padlewski, Daniel Salz, Sebastian Goodman, Adam Grycner, Basil Mustafa, Lucas Beyer, et al. 2022. PaLI: A jointly-scaled multilingual language-image model. arXiv:2209.06794.
- Collomosse and Parsons (2024) John Collomosse and Andy Parsons. 2024. To Authenticity, and Beyond! Building safe and fair generative AI upon the three pillars of provenance. IEEE Computer Graphics and Applications 44, 3 (2024), 82–90.
- Cortes and Vapnik (1995) Corinna Cortes and Vladimir Naumovich Vapnik. 1995. Support-Vector Networks. Machine Learning 20 (1995), 273–297.
- Cozzolino et al. (2024) Davide Cozzolino, Giovanni Poggi, Riccardo Corvi, Matthias Nießner, and Luisa Verdoliva. 2024. Raising the Bar of AI-generated Image Detection with CLIP. In IEEE/CVF Conference on Computer Vision and Pattern Recognition. 4356–4366.
- Datta et al. (2024) Soumyya Kanti Datta, Shan Jia, and Siwei Lyu. 2024. Exposing Lip-syncing Deepfakes from Mouth Inconsistencies. arXiv:2401.10113.
- De Rosa et al. (2024) Vincenzo De Rosa, Fabrizio Guillaro, Giovanni Poggi, Davide Cozzolino, and Luisa Verdoliva. 2024. Exploring the Adversarial Robustness of CLIP for AI-generated Image Detection. arXiv:2407.19553.
- Debarba et al. (2020) Henrique Galvan Debarba, Sylvain Chagué, and Caecilia Charbonnier. 2020. On the plausibility of virtual body animation features in virtual reality. IEEE Transactions on Visualization and Computer Graphics 28, 4 (2020), 1880–1893.
- Diel et al. (2021) Alexander Diel, Sarah Weigelt, and Karl F Macdorman. 2021. A meta-analysis of the uncanny valley’s independent and dependent variables. ACM Transactions on Human-Robot Interaction 11, 1 (2021), 1–33.
- Farid (2022) Hany Farid. 2022. Creating, using, misusing, and detecting deep fakes. Journal of Online Trust and Safety 1, 4 (2022).
- Gen3 ([n. d.]) RunwayML Gen3. [n. d.]. https://runwayml.com/research/introducing-gen-3-alpha.
- Jia et al. (2022) Shan Jia, Xin Li, and Siwei Lyu. 2022. Model attribution of face-swap deepfake videos. In IEEE International Conference on Image Processing. IEEE, 2356–2360.
- Jia et al. (2024) Shan Jia, Reilin Lyu, Kangran Zhao, Yize Chen, Zhiyuan Yan, Yan Ju, Chuanbo Hu, Xin Li, Baoyuan Wu, and Siwei Lyu. 2024. Can ChatGPT detect deepfakes? A study of using multimodal large language models for media forensics. In International Conference on Computer Vision and Pattern Recognition Workshop. 4324–4333.
- Khan and Dang-Nguyen (2024) Sohail Ahmed Khan and Duc-Tien Dang-Nguyen. 2024. CLIPping the Deception: Adapting Vision-Language Models for Universal Deepfake Detection. In International Conference on Multimedia Retrieval. 1006–1015.
- Kondratyuk et al. (2023) Dan Kondratyuk, Lijun Yu, Xiuye Gu, José Lezama, Jonathan Huang, Rachel Hornung, Hartwig Adam, Hassan Akbari, Yair Alon, Vighnesh Birodkar, et al. 2023. VideoPoet: A large language model for zero-shot video generation. arXiv:2312.14125.
- Koukounas et al. (2024) Andreas Koukounas, Georgios Mastrapas, Michael Günther, Bo Wang, Scott Martens, Isabelle Mohr, Saba Sturua, Mohammad Kalim Akram, Joan Fontanals Martínez, Saahil Ognawala, et al. 2024. Jina CLIP: Your CLIP Model Is Also Your Text Retriever. arXiv:2405.20204.
- Lin and Yang (2024) Shanchuan Lin and Xiao Yang. 2024. AnimateDiff-Lightning: Cross-model diffusion distillation. arXiv:2403.12706.
- Loper et al. (2023) Matthew Loper, Naureen Mahmood, Javier Romero, Gerard Pons-Moll, and Michael J Black. 2023. SMPL: A skinned multi-person linear model. In Seminal Graphics Papers: Pushing the Boundaries, Volume 2. 851–866.
- McDonnell et al. (2012) Rachel McDonnell, Martin Breidt, and Heinrich H Bülthoff. 2012. Render me real? Investigating the effect of render style on the perception of animated virtual humans. ACM Transactions on Graphics 31, 4 (2012), 1–11.
- Neri et al. (1998) Peter Neri, M Concetta Morrone, and David C Burr. 1998. Seeing biological motion. Nature 395, 6705 (1998), 894–896.
- Ng et al. (2024) Evonne Ng, Javier Romero, Timur Bagautdinov, Shaojie Bai, Trevor Darrell, Angjoo Kanazawa, and Alexander Richard. 2024. From audio to photoreal embodiment: Synthesizing humans in conversations. In IEEE/CVF Conference on Computer Vision and Pattern Recognition. 1001–1010.
- Nightingale and Farid (2022) Sophie J Nightingale and Hany Farid. 2022. AI-synthesized faces are indistinguishable from real faces and more trustworthy. Proceedings of the National Academy of Sciences 119, 8 (2022), e2120481119.
- Nirkin et al. (2019) Yuval Nirkin, Yosi Keller, and Tal Hassner. 2019. FSGAN: Subject agnostic face swapping and reenactment. In IEEE/CVF International Conference on Computer Vision. 7184–7193.
- Nirkin et al. (2021) Yuval Nirkin, Lior Wolf, Yosi Keller, and Tal Hassner. 2021. Deepfake detection based on discrepancies between faces and their context. IEEE Transactions on Pattern Analysis and Machine Intelligence 44, 10 (2021), 6111–6121.
- Pexels ([n. d.]) Pexels. [n. d.]. https://www.pexels.com.
- Pianese et al. (2022) Alessandro Pianese, Davide Cozzolino, Giovanni Poggi, and Luisa Verdoliva. 2022. Deepfake audio detection by speaker verification. In IEEE International Workshop on Information Forensics and Security. IEEE, 1–6.
- Radford et al. (2021) Alec Radford, Jong Wook Kim, Chris Hallacy, Aditya Ramesh, Gabriel Goh, Sandhini Agarwal, Girish Sastry, Amanda Askell, Pamela Mishkin, Jack Clark, et al. 2021. Learning transferable visual models from natural language supervision. In International Conference on Machine Learning. PMLR, 8748–8763.
- Schuhmann et al. (2022) Christoph Schuhmann, Romain Beaumont, Richard Vencu, Cade Gordon, Ross Wightman, Mehdi Cherti, Theo Coombes, Aarush Katta, Clayton Mullis, Mitchell Wortsman, et al. 2022. LAION-5B: An open large-scale dataset for training next generation image-text models. Advances in Neural Information Processing Systems 35 (2022), 25278–25294.
- Schuhmann et al. (2021) Christoph Schuhmann, Richard Vencu, Romain Beaumont, Robert Kaczmarczyk, Clayton Mullis, Aarush Katta, Theo Coombes, Jenia Jitsev, and Aran Komatsuzaki. 2021. LAION-400M: Open dataset of clip-filtered 400 million image-text pairs. arXiv:2111.02114.
- Shen et al. (2021) Sheng Shen, Liunian Harold Li, Hao Tan, Mohit Bansal, Anna Rohrbach, Kai-Wei Chang, Zhewei Yao, and Kurt Keutzer. 2021. How much can CLIP benefit vision-and-language tasks? arXiv:2107.06383.
- Smeu et al. (2024) Stefan Smeu, Elisabeta Oneata, and Dan Oneata. 2024. DeCLIP: Decoding CLIP representations for deepfake localization. arXiv:2409.08849.
- Song et al. (2022) Haoyu Song, Li Dong, Wei-Nan Zhang, Ting Liu, and Furu Wei. 2022. CLIP models are few-shot learners: Empirical studies on vqa and visual entailment. arXiv:2203.07190.
- Sora ([n. d.]) Sora. [n. d.]. https://openai.com/index/sora/.
- Suwajanakorn et al. (2017) Supasorn Suwajanakorn, Steven M. Seitz, and Ira Kemelmacher-Shlizerman. 2017. Synthesizing Obama: Learning lip sync from audio. ACM Transactions on Graphics (2017).
- Vahdati et al. (2024) Danial Samadi Vahdati, Tai D Nguyen, Aref Azizpour, and Matthew C Stamm. 2024. Beyond Deepfake Images: Detecting AI-Generated Videos. In IEEE/CVF Conference on Computer Vision and Pattern Recognition, Workshop on Media Forensics. 4397–4408.
- Veo ([n. d.]) Veo. [n. d.]. https://deepmind.google/technologies/veo.
- Voloshynovskiy et al. (2001) Sviatoslav Voloshynovskiy, Shelby Pereira, Thierry Pun, Joachim J Eggers, and Jonathan K Su. 2001. Attacks on digital watermarks: Classification, estimation based attacks, and benchmarks. IEEE Communications Magazine 39, 8 (2001), 118–126.
- Yang et al. (2024) Zhuoyi Yang, Jiayan Teng, Wendi Zheng, Ming Ding, Shiyu Huang, Jiazheng Xu, Yuanming Yang, Wenyi Hong, Xiaohan Zhang, Guanyu Feng, et al. 2024. CogVideoX: Text-to-video diffusion models with an expert transformer. arXiv:2408.06072.
- Yu et al. (2024) Heng Yu, Chaoyang Wang, Peiye Zhuang, Willi Menapace, Aliaksandr Siarohin, Junli Cao, Laszlo A Jeni, Sergey Tulyakov, and Hsin-Ying Lee. 2024. 4Real: Towards Photorealistic 4D Scene Generation via Video Diffusion Models. arXiv:2406.07472.
- Zhai et al. (2023) Xiaohua Zhai, Basil Mustafa, Alexander Kolesnikov, and Lucas Beyer. 2023. Sigmoid loss for language image pre-training. In IEEE/CVF International Conference on Computer Vision. 11975–11986.

### v2

@@
Abstract
AI-generated video generation continues its journey through the uncanny valley to produce content that is increasingly perceptually indistinguishable from reality. To better protect individuals, organizations, and societies from its malicious applications, we describe an effective and robust technique for distinguishing real from AI-generated human motion using multi-modal semantic embeddings. Our method is robust to the types of laundering that typically confound more low- to mid-level approaches, including resolution and compression attacks. This method is evaluated against DeepAction, a custom-built, open-sourced dataset of video clips with human actions generated by seven text-to-video AI models and matching real footage. The dataset is available under an academic license at https://www.huggingface.co/datasets/faridlab/deepaction_v1.
1 Introduction
@@
Generative AI, unlike earlier model-based animation, has emerged as an intriguing new genre for computer animation. While earlier versions of AI-generated video were things nightmares are made of (see, for example, “Will Smith eating spaghetti”111https://www.youtube.com/watch?v=XQr4Xklqzw8), recent advances have shown significant improvements in photo-realism and temporal consistency.
@@
We have started to see AI-generated images weaponized in the form of child sexual abuse material, non-consensual sexual imagery, fraud, and as an accelerant to disinformation campaigns (Farid, 2022). There is no reason, therefore, to believe that AI-generated video will not follow suit.
@@
Motivated by, and building upon, earlier work, we describe a technique to detect AI-generated videos containing human motion. Our initial focus is on humans because these videos are of most concern when it comes to the harms enumerated above. Despite this focus, we will discuss why our approach is likely to generalize to other types of video content. We evaluate the efficacy of our approach on a diverse dataset of our creation consisting of real and matching (in terms of the human actions depicted) AI-generated videos from seven different generative-AI models. We demonstrate the robustness of our detection in the face of standard laundering attacks like resizing and transcoding, and evaluate its generalizability to previously unseen models. Our work contributes to this nascent literature with the:
@@
1.
@@
5.
construction and release of DeepAction, a new benchmark dataset of real and AI-generated human motion.
2 Related Work
Reactive techniques – operating in the absence of credentials – fall into two basic approaches: (2a) learning-based, in which features that distinguish real from fake content are learned by a range of machine-learning techniques, and (2b) artifact-based, in which a range of low-level (pixel-based) to high-level (semantic-based) features are explicitly extracted to distinguish between real and fake content (Farid, 2022).
Another recent example (Jia et al., 2024) explores the potential of multi-modal large language models (LLMs) for detecting AI-generated faces.The authors prompt ChatGPT with an image and prompt like “Tell me if this is an AI-generated image.” This approach achieves an average accuracy of (as measured by area under the curve, AUC). Although not particularly accurate, what is intriguing about this approach is that it points to a potentially semantic-level reasoning.
Recent studies took a more direct semantic approach by leveraging a contrastive language-image pretraining (CLIP) representation (Radford et al., 2021). In (Cozzolino et al., 2024), the authors extract a CLIP embedding from real and AI-generated images and with only a linear SVM achieve detection accuracy ranging from to depending on the amount of image post-processing.
3 Dataset
We generated video clips from seven text-to-video AI models: BD AnimateDiff (Lin and Yang, 2024), CogVideoX-5B (Yang et al., 2024), Lumiere222An earlier version of the Lumiere model was used which was not specifically trained or fine-tuned on human motion. (Bar-Tal et al., 2024), RunwayML Gen3 (Gen3, 2024), Stable Diffusion Txt2Img+Img2Vid (Blattmann et al., 2023), Veo333A pre-release version of the Veo model was used. (Veo, 2024), and VideoPoet (Kondratyuk et al., 2023). The default generation parameters were used for each model. These AI-generated videos are minutes in length constituting video frames, Figure 2.
@@ 3 Dataset
We also curated a set of real videos from Pexels (Pexels, 2024), an open-source stock video database, matched to the human-action prompts described above. These real videos are minutes in length constituting individual video frames. These videos were matched on general action so that the embedding representations (described next) between real and fake would not be based on semantic differences.
The resulting dataset, called DeepAction, comprises eight video categories (seven AI-generated and one real) and is made publicly available.
4 Methods
We describe four CLIP-like, multi-modal embeddings and proceed with classification schemes used to distinguish the real from the fake.
4.1 Embeddings
@@
1.
The CLIP embedding model was trained on LAION-2B (Schuhmann et al., 2022). We used the smallest -D baseline model ViT-B/32.
2.
The SigLIP embedding model was trained on WebLI (Chen et al., 2022), which is over larger than LAION-2B. Unlike CLIP’s softmax-based contrastive loss, SigLIP uses a sigmoid loss. We used the -D patch16-224 model variant.
3.
The JinaCLIP embedding model was trained on LAION-400M (Schuhmann et al., 2021) along with an additional set of text-pair datasets. While CLIP’s training optimized for text-image representation alignment, JinaCLIP was trained to jointly optimize text-image and text-text representation alignment. We used the base -D v1 model.
4.
@@
See Appendix for baseline model and training details.
4.2 Classification
@@
1.
A two-class SVM (Cortes and Vapnik, 1995), implemented in scikit-learn444https://www.scikit-learn.org/stable/modules/generated/sklearn.svm.SVC.html, is used to classify video frames as real or fake, in which all seven text-to-video models are bundled into a single ’fake’ class.
2.
@@ 5 Results
For the CLIP, SigLIP, and JinaCLIP embeddings, our dataset is randomly split into an train/test partition. For the fine-tuned CLIP, the dataset is randomly split into a partition, where the first is used for CLIP fine-tuning, is used for model training, and is used for testing. The splits are determined at the action (prompt) level to ensure no overlap across partitions.
In each case, we report the mean frame- and video-level accuracy on the test set, averaged over five random train/test repetitions. Because our dataset is imbalanced, with significantly more fake than real videos, we under-sample the fake videos. Throughout, we report accuracy as a macro-average by evenly weighting the class accuracies.
@@ 5.1 Robustness
Shown in the left portion of Table 3 are the frame- and video-level accuracies for the CLIP embedding and a two-class linear SVM for videos with progressively lower resolution (we quantify the change in resolution as a percentage because the original videos were of varying resolution, Section 3). At the video level, accuracy remains relatively high for videos at or higher of their original resolution. This corresponds to an average resolution of .
@@ 5.2 Generalizability
5.3 Non-Human Motion
Having been trained on only videos containing human motion, we wondered if our two-class SVM would generalize to arbitrary text-to-video models. We evaluated our model on videos generated by Sora (Sora, 2024) and RunwayML (Gen3, 2024) that did not contain any humans or human-motion. With the CLIP embeddings, the model correctly classified of these videos as AI-generated showing generalization to non-human motion and to one text-to-video model not seen in training (Sora).
@@ 5.5 CGI
Performance for SigLIP across all prompts was perfect at . That is, the SigLIP embedding for these video frames is more similar to the “AI-generated”, “unrealistic”, or “manipulated” prompts than the “authentic” or “realistic” prompts. We don’t fully understand why there is such a large discrepancy here across embedding, but it may be possible that SigLIP was exposed to CGI content during training.
5.6 Comparison to Related Work
Due to the lack of standardized benchmarks, comparing methods remains imperfect. Nevertheless, we provide a comparison to related work by comparing baseline accuracy and generalizability to unseen generative models. The latter is particularly important because true efficacy in the wild will be limited by the ability of detection models to generalize.
@@ 5.6 Comparison to Related Work
6 Discussion
When we first started to think about the problem of distinguishing real from AI-generated human motion, we thought to take a physics-based approach leveraging advances in 3D human-body modeling (Loper et al., 2023). We hypothesized that by extracting 3D models of the human body, we could expose implausible dynamics and kinematics in AI-generated motion. We, however, quickly ran into obstacles and found it difficult to consistently and reliably extract 3D models in a wide range of human poses and levels of occlusion.
@@ 6 Discussion
Ethical Statement
Unlike the ethical considerations that should be weighed in the creation of generative-AI models (e.g., training data ownership and potential misuse in the form of the creation of non-consensual intimate imagery), we don’t anticipate similar concerns with the development of techniques to detect AI-generated content. We recognize, however, that describing a detection technique could further enable the creation of more compelling deepfakes. We believe, however, that the benefit to the scientific field outweighs this risk, and to mitigate this risk we have chosen not to open-source our code.
Acknowledgements
@@
This work was done during the first author’s (MB) internship at Google. The first author thanks his manager, Tony Chen, and co-host, Catherine Zhao, for their exceptional support, mentoring, and discussions related to this work; thanks to Prakash Hariramani, Sunita Verma, and Joshy Joseph for their support within leadership; and thanks for helpful discussions to: Amanda Walker, Athula Balachandran, Dara Bahri, David Hendon, Jess Gallegos, John Choi, Karen Guo, Nicholas Carlini, Nick Pezzoti, Stanley Wang, Sven Gowal, and Tom Hume. This work was supported by a gift from YouTube and funding from the University of California Noyce Initiative.
Appendix A Experimental Setup
A.1 CLIP
The following model was used:
A.2 SigLIP
The following model was used:
A.3 JinaCLIP
The following model was used:
A.4 FT-CLIP
The model was fine-tuned for one warm-up epoch at a learning rate of and one regular epoch at a learning rate of . An SGD optimizer with a cosine learning rate scheduler was employed for both epochs. The batch size was set to for training and for testing. The training set comprised two balanced sets of real and AI-generated frames, with captions constructed as ”a {REAL/FAKE} image of {ACTION PROMPT}”. Training was performed on an A100 GPU for approximately eight hours.